OpenStudy (anonymous):

Which is the solution set of -4 <= 3- y over the whole numbers? {0, 1, 2, 3, 4, 5, 6} {6, 7, 8, 9, 10, …} {7, 8, 9, 10, …} {0, 1, 2, 3, 4, 5, 6, 7}

jimthompson5910 (jim_thompson5910):

First I'd add y to both sides -4 <= 3- y -4+y <= 3- y+y y + (-4) <= 3 y - 4 <= 3 What is the next step to isolate y on the left side?

OpenStudy (anonymous):

add 4 to both sides right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

resulting in what?

OpenStudy (anonymous):

y<7 is that right?

jimthompson5910 (jim_thompson5910):

y <= 7 actually, but you have the right idea

jimthompson5910 (jim_thompson5910):

which set of whole numbers represents y <= 7

OpenStudy (anonymous):

okay and I'm not sure /: how can i tell?

jimthompson5910 (jim_thompson5910):

what numbers make y <= 7 true?

jimthompson5910 (jim_thompson5910):

does y = 10 make y <= 7 true?

OpenStudy (anonymous):

no it doesnt :)

jimthompson5910 (jim_thompson5910):

how about y = 9?

OpenStudy (anonymous):

nope :)

jimthompson5910 (jim_thompson5910):

so basically the set of numbers less than 7 (7 is included) makes y <= 7 true numbers like 0, 2, or 6 are part of this set

jimthompson5910 (jim_thompson5910):

each answer choice is simply listing out each whole number solution

OpenStudy (anonymous):

so the last set is the answer ?

jimthompson5910 (jim_thompson5910):

yep {0, 1, 2, 3, 4, 5, 6, 7}

OpenStudy (anonymous):

Oh thank you ! :)

jimthompson5910 (jim_thompson5910):

you can also test those values in -4 <= 3- y and it should be true as well